Lear Inc. has $1,000,000 in current assets, $450,000 of which are considered permanent current assets. In addition, the firm has $800,000 invested in fixed assets.    
  
a. Lear wishes to finance all fixed assets and half of its permanent current assets with long-term financing costing 9 percent. The balance will be financed with short-term financing, which currently costs 7 percent. Lear’s earnings before interest and taxes are $400,000. Determine Lear’s earnings after taxes under this financing plan. The tax rate is 30 percent.
  

Earning After Taxes

b. As an alternative, Lear might wish to finance all fixed assets and permanent current assets plus half of its temporary current assets with long-term financing and the balance with short-term financing. The same interest rates apply as in part a. Earnings before interest and taxes will be $400,000. What will be Lear’s earnings after taxes? The tax rate is 30 percent.

Earnings After Taxes
Answer #1

a. Short-term interest expense = $ ( 550,000 + 225,000) x 7% = $ 54,250

Long-term interest expense = $ ( 800,000 + 225,000) x 9 % = $ 92,250

Total Interest Expense = $ 54,250 + $ 92,250 = $ 146,500

Earnings before taxes = $ 400,000 - $ 146,500 = $ 253,500

Earnings after taxes = $ 253,500 x ( 1 - 0.30) = $ 177,450.

b. Long-term interest expense = $ ( 800,000 + 450,000 + 275,000 ) x 9 % = $ 137,250

Short-term interest expense = $ 275,000 x 7 % = $ 19,250

Total interest expense = $ 137,250 + $ 19,250 = $ 156,500.

Earnings before taxes = $ 400,000 - $ 156,500 = $ 243,500.

Earnings after taxes = $ 243,500 x ( 1 - 0.30 ) = $ 170,450.